# Precision gravity constraints on large dark sectors

### v1

@@ Abstract
General relativity, treated as a low-energy effective field theory, predicts quantum corrections to Newton’s law of gravitation arising from loops of matter and graviton fields. While these corrections are utterly negligible when considering the Standard Model particle content, the situation changes dramatically in the presence of a hidden or dark sector containing a very large number of light degrees of freedom. In such cases, loop-induced modifications to the Newtonian potential can accumulate to levels testable in laboratory and astrophysical probes of gravity at short distances. In this work we systematically derive and constrain the impact of large dark sectors on precision tests of Newton’s law, translating effective field theory predictions into the experimental language of Yukawa-type deviations and inverse-square law deformations. By mapping precision fifth–force constraints onto bounds on species multiplicities and masses, we show that current and future experiments already impose nontrivial constraints on the size and structure of hidden sectors coupled only gravitationally. Our results provide a model-independent framework for confronting dark-sector scenarios with precision gravity data, and highlight the complementarity of particle physics and short-distance gravitational tests in probing new physics beyond the Standard Model.
@@ 1 Introduction
Einstein’s general relativity (GR) continues to be the most successful classical description of gravity, having passed all experimental tests from the Solar System to binary pulsars with remarkable accuracy. However, as is by now well recognized, GR should be regarded as a low-energy effective field theory (EFT), valid below the Planck scale $M_{\rm Pl}$, where quantum effects become non-negligible Donoghue (1994). Within this EFT framework, quantum loops of matter and gravitons induce calculable corrections to the Newtonian potential, with leading terms scaling as $1/r^{3}$ at large distances Bjerrum-Bohr et al. (2003); Holstein (2023); Faller (2008). The formalism of covariant nonlocal form factors, developed through the Barvinsky–Vilkovisky heat-kernel expansion, provides a powerful language to encode such quantum corrections Codello and Jain (2015, 2016).
@@ 2.2 Effective Action and Nonlocal Form Factors
The quantum corrections to gravity can be systematically studied within the framework of the one-loop effective action, treating general relativity as a low-energy effective field theory (EFT). For a matter field of mass $m$ and spin $s$, the one-loop effective action takes the generic form Codello and Jain (2016, 2015)
$\Gamma^{(1)}_{s}\;=\;\frac{1}{2}(-1)^{2s}\,\mathrm{Tr}\,\ln\big{(}\Delta_{s}+m^{2}\big{)},$ (1)
@@
where $\Delta_{s}$ is the kinetic operator appropriate to spin $s$. The nonlocal terms of interest appear as form factors multiplying curvature invariants,
$\Gamma_{\text{eff}}\supset\int d^{4}x\sqrt{g}\,R\,F_{s}(-\Box/m^{2})\,R+\ldots,$ (2)
where $F_{s}$ encodes the nonlocal (scale-dependent) quantum correction Codello and Jain (2016, 2015). The presence of $F_{s}$ reflects the fact that quantum fluctuations generate momentum-dependent corrections rather than simply renormalizing Newton’s constant.
For arbitrary spin $s$ (real scalar, Dirac fermion, and vector), the Barvinsky–Vilkovisky/heat-kernel expansion yields Codello and Jain (2016); Fröb (2016)
$F_{s}(\tau)\;=\;c_{s}\left[\,1-\tau\ln\!\left(1+\frac{1}{\tau}\right)\,\right],\qquad\tau\equiv-\Box/m^{2},$ (3)
with spin-dependent coefficients
@@
$c_{0}$ $=\frac{1}{384\pi^{2}},$ $\text{(real scalar)},$ (4)
$c_{1/2}$ $=\frac{1}{192\pi^{2}},$ $\text{(Dirac fermion)},$ (5)
$c_{1}$ $=-\frac{1}{96\pi^{2}},$ $\text{(vector, with ghosts included)}.$ (6)
These constants determine the relative size and sign of the nonlocal corrections, and are ultimately tied to the number of dynamical degrees of freedom carried by the particle.
The nonlocal term above leads to $1/r^{3}$-type corrections in the gravitational potential. For example, integrating out a real scalar yields a correction
@@
If the loop particle is massless, $\tau\to\infty$, and the form factors reduce to the pure nonanalytic structure $F_{s}(\tau)\sim c_{s}\ln(-\Box/\mu^{2})$, with $\mu$ a renormalization scale. This directly implies a scale-dependent gravitational coupling $G(q^{2})$. We discuss the massless case in more detail in the following section. For particles with nonzero mass, the mass scale $m$ acts as a threshold: at distances much larger than $1/m$, the quantum corrections are strongly suppressed, while at shorter distances the corrections gradually approach the same logarithmic running seen in the massless case. In this way, the massive case smoothly connects between $1/m^{2}$-suppressed effects in the low energy ($q^{2}\ll m^{2}$) and universal logarithmic behavior in the high energy case ($q^{2}\gg m^{2}$), a feature that plays an important role in precision tests of gravity across laboratory and astrophysical regimes.
@@ 2.3 Running and Threshold Effects in $G(q^{2})$
where $\mu$ is a renormalization scale. This result is explicitly derived and discussed in detail later in Section 3. This universal logarithmic dependence is responsible for long-distance, nonlocal corrections to Newton’s law Donoghue (1994); Bjerrum-Bohr et al. (2003); Kirilin and Khriplovich (2002).
@@ 2.4 Correction to the Newtonian Potential
As outlined above, the corrections to the Newtonian potential are obtained by integrating the nonlocal form factor above for all possible particle spins $s$. In the case of the massless fields, such corrections to the gravitational potential induce long range effects of the form $O(1/r^{3})$ Donoghue (1994); Bjerrum-Bohr et al. (2003); Kirilin and Khriplovich (2002); Faller (2008); Holstein (2023):
@@ 3.1 Momentum-Space Running
For a field of mass $m$ and spin $s$, the nonlocal correction to Newton’s constant in momentum space is obtained via integrating the quantum corrections given by $F_{s}$ in Eq. 3 to obtain the form  Fröb (2016):
@@
where $\mu$ has replaced the mass $m$ and defines the renormalization scale in the massless case. This matches the standard massless result, where the running of $G$ is governed by a $q^{2}\ln(-q^{2})$ form factor Donoghue (1994); Bjerrum-Bohr et al. (2003). In the limit $m\rightarrow 0$, the nonlocal form factor reduces to the universal massless expression $\sim q^{2}ln(-q^{2})$, in agreement with earlier EFT analyses of graviton and massless matter loops. The scale m in the massive case thus plays the role of a regulator, which is replaced by a renormalization scale $\mu$ in the massless theory.
@@ 5.3 Massive dark degrees of freedom: mapping to Yukawa envelopes
Setting $\mathcal{S}_{X}(1/2)=1$ gives a slightly conservative estimate that is convenient for quick re-interpretations of published $(\alpha,\lambda)$ curves (torsion balances, LLR, planetary dynamics, etc.). Applying the results from precision experiments at varying length scales, we can obtain continuous limits on the multiplicities of dark sectors, as is illustrated in the figure below.
Figure 1: Degrees of freedom set by experimental constrains on $G$ and expansion parameters $\alpha$ and $\lambda$. Shaded region is exlusion zone set by experimental limits at each distance scale $r_{*}$.
@@ 6 Discussion and Conclusions
Our results have direct implications for several classes of theoretical models. Strongly-coupled dark gauge theories can produce large numbers of light mesons and baryons, and while individual hadrons contribute negligibly, a rich spectrum with hundreds or thousands of states could collectively generate detectable gravitational effects. Our bounds provide quantitative guidance for the maximum complexity such sectors can achieve while remaining consistent with laboratory tests. Extra-dimensional theories often predict towers of Kaluza-Klein modes or large numbers of moduli fields, and the gravitational constraints complement existing bounds from fifth-force searches to help restrict the allowed parameter space of phenomenologically viable compactifications. Similar analyses could be extended to axion-like particles that couple to gravity through topological terms or induce effective modifications to gravitational couplings.

### v2

@@ Abstract
General relativity, treated as a low-energy effective field theory, predicts quantum corrections to Newton’s law of gravitation arising from loops of matter and graviton fields. While these corrections are utterly negligible for the Standard Model particle content, the situation changes dramatically in the presence of a hidden or dark sector containing a very large number of light degrees of freedom. In such cases, loop-induced modifications to the Newtonian potential can accumulate to levels testable in laboratory and astrophysical probes of gravity at short distances. In this work we systematically derive and constrain the impact of large dark sectors on precision tests of Newton’s law, translating effective-field-theory predictions into the experimental language of Yukawa-type deviations and inverse-square-law deformations. By mapping precision fifth–force constraints onto bounds on species multiplicities and masses, we show that current and forthcoming experiments already impose nontrivial constraints on the size and structure of hidden sectors coupled only gravitationally. For truly massless hidden states, present data still permit multiplicities as large as $\mathcal{O}(10^{61})$, with modest spin dependence; for finite masses the constraints reduce to the familiar short-range Yukawa parameterization. Our results provide a model-independent framework for confronting dark-sector scenarios with precision gravity data and clarify how non-minimal scalar couplings, potential higher-derivative poles at large species number, and Kaluza–Klein towers fit within this picture. The approach is complementary to cosmological probes: Big Bang Nucleosynthesis and the Cosmic Microwave Background constrain relic abundances under specified production histories, whereas laboratory tests constrain the spectrum of light states irrespective of their cosmological population.
@@ 1 Introduction
Einstein’s general relativity (GR) continues to be the most successful classical description of gravity, having passed all experimental tests from the Solar System to binary pulsars with remarkable accuracy. However, as is by now well recognized, GR should be regarded as a low-energy effective field theory (EFT), valid below the Planck scale $M_{\rm Pl}$, where quantum effects become non-negligible Donoghue (1994a). Within this EFT framework, quantum loops of matter and gravitons induce calculable corrections to the Newtonian potential, with leading terms scaling as $1/r^{3}$ at large distances Bjerrum-Bohr et al. (2003); Holstein (2023); Faller (2008). The formalism of covariant nonlocal form factors, developed through the Barvinsky–Vilkovisky heat-kernel expansion, provides a powerful language to encode such quantum corrections Codello and Jain (2015, 2016).
@@ 2.2 Effective Action and Nonlocal Form Factors
Quantum corrections to gravity can be systematically treated within the one-loop effective action, viewing general relativity as a low-energy EFT. For a matter field of mass $m$ and spin $s$ one writes Codello and Jain (2016, 2015)
$\Gamma^{(1)}_{s}\;=\;\tfrac{1}{2}(-1)^{2s}\,\mathrm{Tr}\,\ln\!\big(\Delta_{s}+m^{2}\big),$ (1)
@@
where $\Delta_{s}$ is the kinetic operator appropriate to spin $s$. The nonlocal terms of interest appear as form factors multiplying curvature invariants. It is convenient and complete to organize the result in the quadratic curvature basis (see, e.g., Barvinsky and Vilkovisky (1990); Donoghue (1994b); Donoghue and Holstein (2017); Burns and Pilaftsis (2015); Fröb et al. (2015)):
$\Gamma^{(2)}_{\rm matter}=\int d^{4}x\,\sqrt{-g}\Big[R\,F_{1,s}\!\Big(\tfrac{\Box}{m^{2}}\Big)R+R_{\mu\nu}\,F_{2,s}\!\Big(\tfrac{\Box}{m^{2}}\Big)R^{\mu\nu}\Big],$ (2)
with the dimensionless argument
$\tau\;\equiv\;-\,\frac{\Box}{m^{2}}.$ (3)
To leading nonlocal order, each spin-$s$ species contributes the universal shape
$F_{i,s}(\tau)=c_{i,s}\Big[\,1-\tau\ln\!\Big(1+\frac{1}{\tau}\Big)\Big],\qquad i=1,2,$ (4)
which resums the infrared logarithms and smoothly interpolates between the massive and massless regimes.
Note that in the background–field quantization of a (Stueckelberg-regulated) Abelian vector in a covariant gauge, the one-loop effective action at quadratic order reads schematically
$\Gamma^{(1)}_{(1)}\;=\;\tfrac{1}{2}\,\mathrm{Tr}\ln\!\big(\Delta_{\mu}{}^{\nu}+m^{2}\big)\;-\;\mathrm{Tr}\ln\!\big(\Delta_{\rm gh}+m^{2}\big)\;+\;\tfrac{1}{2}\,\mathrm{Tr}\ln\!\big(\Delta_{\rm gf}+m^{2}\big),$
where the first term comes from the gauge field, the second from the complex Grassmann Faddeev–Popov ghosts (hence the overall minus sign), and the last from the gauge-fixing Jacobian. Decomposing into spin projectors and rewriting the result in the curvature basis gives
$\Gamma^{(2)}_{\rm matter}\supset\!\int\!\sqrt{-g}\,\Big[R\,F_{1,1}\!\Big(\tfrac{\Box}{m^{2}}\Big)R\;+\;R_{\mu\nu}\,F_{2,1}\!\Big(\tfrac{\Box}{m^{2}}\Big)R^{\mu\nu}\Big],$
with the same nonlocal shape for both channels, $F_{i,1}(\tau)=c_{i,1}\,[\,1-\tau\ln(1+1/\tau)\,]$, but with spin-1 coefficients $c_{i,1}$ that include the ghost/gauge-fixing contributions. The static inverse–square–law combination probed by our fits is $F^{\rm(ISL)}_{1}(\tau)=F_{2,1}(\tau)-\tfrac{1}{2}F_{1,1}(\tau)$, so the net coefficient is $c_{1}\equiv c_{2,1}-\tfrac{1}{2}c_{1,1}.$ Including the Faddeev–Popov sector yields a negative value for this combination, $c_{1}=-1/(96\pi^{2})$; see, e.g., Barvinsky and Vilkovisky (1990); Burns and Pilaftsis (2015); Fröb et al. (2015) for explicit derivations in the covariant form-factor language.
At the linearized level about flat space, the static potential probes a definite combination of the spin-2 and spin-0 channels carried by the propagator. This can be expressed as
$F^{\rm(ISL)}_{s}(\tau)=F_{2,s}(\tau)-\tfrac{1}{2}F_{1,s}(\tau)=c_{s}\Big[\,1-\tau\ln\!\Big(1+\frac{1}{\tau}\Big)\Big],\qquad c_{s}\equiv c_{2,s}-\tfrac{1}{2}c_{1,s}.$ (5)
Equation (5) shows explicitly that both $R^{2}$ and $R_{\mu\nu}R^{\mu\nu}$ are generated at one loop, while clarifying why a single coefficient $c_{s}$ suffices for the static potential: it is the short-distance combination entering inverse–square–law (ISL) observables.
For minimally coupled matter fields, the coefficients entering the ISL combination are
@@
$c_{0}=\frac{1}{384\,\pi^{2}}\ \ \text{(real scalar)},\quad c_{1/2}=\frac{1}{192\,\pi^{2}}\ \ \text{(Dirac fermion)},\quad c_{1}=-\,\frac{1}{96\,\pi^{2}}\ \ \text{(vector+ghosts)}.$ (6)
Summing over all species gives
$F^{\rm(ISL)}(\tau)\;=\;\sum_{s}N_{s}\,c_{s}\,\Big[\,1-\tau\ln\!\Big(1+\frac{1}{\tau}\Big)\Big],$ (7)
where $N_{s}$ counts the number of fields of spin $s$. If a non-minimal scalar coupling $\tfrac{1}{2}\xi R\phi^{2}$ is present, the trace (spin-0) channel carries the familiar $(1-6\xi)^{2}$ weight Barvinsky and Vilkovisky (1990); Burns and Pilaftsis (2015); in practice one may track this by $c_{0}\!\to c_{0}(1-6\xi)^{2}$ in (7).
In the massless limit ($\tau\to\infty$), $F_{i,s}(\tau)\simeq c_{i,s}(\ln\tau-2)$, reproducing the standard running $\ln(-\Box/\mu^{2})$ Barvinsky and Vilkovisky (1990); Donoghue (1994b). For $|\tau|\ll 1$ (momenta well below $m$),
$F_{i,s}(\tau)=c_{i,s}\Big(\frac{\tau}{6}-\frac{\tau^{2}}{60}+\cdots\Big),$ (8)
which yields the local derivative series generating the universal long-distance $1/r^{3}$ tail in the static potential. At intermediate distances the same physics is conveniently encoded by a Yukawa-envelope fit Fröb et al. (2015); Burns and Pilaftsis (2015). In this way, our mapping from ISL data to bounds on the cumulative contribution of hidden fields naturally accounts for both curvature–squared operators and their nonlocal running.
For example, integrating out a real scalar yields a correction
@@
If the loop particle is massless, thus $\tau\to\infty$, the form factors reduce to the pure nonanalytic structure $F_{s}(\tau)\sim c_{s}\ln(-\Box/\mu^{2})$, with $\mu$ a renormalization scale. This directly implies a scale-dependent gravitational coupling $G(q^{2})$. We discuss the massless case in more detail in the following section. For particles with nonzero mass, the mass scale $m$ acts as a threshold: at distances much larger than $1/m$, the quantum corrections are strongly suppressed, while at shorter distances the corrections gradually approach the same logarithmic running seen in the massless case. In this way, the massive case smoothly connects between $1/m^{2}$-suppressed effects in the low energy ($q^{2}\ll m^{2}$) and universal logarithmic behavior in the high energy case ($q^{2}\gg m^{2}$), a feature that plays an important role in precision tests of gravity across laboratory and astrophysical regimes.
@@ 2.3 Running and Threshold Effects in $G(q^{2})$
where $\mu$ is a renormalization scale. This result is explicitly derived and discussed in detail later in Section 3. This universal logarithmic dependence is responsible for long-distance, nonlocal corrections to Newton’s law Donoghue (1994a); Bjerrum-Bohr et al. (2003); Kirilin and Khriplovich (2002).
@@ 2.4 Correction to the Newtonian Potential
As outlined above, the corrections to the Newtonian potential are obtained by integrating the nonlocal form factor above for all possible particle spins $s$. In the case of the massless fields, such corrections to the gravitational potential induce long range effects of the form $O(1/r^{3})$ Donoghue (1994a); Bjerrum-Bohr et al. (2003); Kirilin and Khriplovich (2002); Faller (2008); Holstein (2023):
@@
2.5 Validity of the EFT and species effects.
Loop-generated quadratic curvature terms can be reorganized into massive spin-0 and spin-2 propagating poles (the latter being a ghost in a purely local $R^{2}\!+\!R_{\mu\nu}R^{\mu\nu}$ truncation). In a Wilsonian picture the induced masses scale schematically like
$M_{0,2}\;\sim\;\frac{M_{\rm Pl}}{\sqrt{\alpha_{0,2}}}\,,\qquad\alpha_{0,2}\;\propto\;\frac{N}{16\pi^{2}},$
[…]
@@ 3.1 Momentum-Space Running
For a field of mass $m$ and spin $s$, the nonlocal correction to Newton’s constant in momentum space is obtained via integrating the quantum corrections given by $F_{s}$ in Eq. LABEL:eq:quantum_corr to obtain the form  Fröb (2016):
@@
where $\mu$ has replaced the mass $m$ and defines the renormalization scale in the massless case. This matches the standard massless result, where the running of $G$ is governed by a $q^{2}\ln(-q^{2})$ form factor Donoghue (1994a); Bjerrum-Bohr et al. (2003). In the limit $m\rightarrow 0$, the nonlocal form factor reduces to the universal massless expression $\sim q^{2}ln(-q^{2})$, in agreement with earlier EFT analyses of graviton and massless matter loops. The scale m in the massive case thus plays the role of a regulator, which is replaced by a renormalization scale $\mu$ in the massless theory.
@@ 4 Constraints on Hidden or Dark Sectors with Many Degrees of Freedom
Note that our reinterpretation counts any propagating degree of freedom lighter than the experimental lever arm. Concretely, a state contributes to the potential if its mass satisfies $m\lesssim r_{\star}^{-1}$. In confining theories, heavy resonances above this scale are integrated out and do not affect our fits, whereas light PNGBs, glueballs, or other composites couple universally to gravity and contribute on the same footing as elementary fields. The bounds therefore depend only on the infrared spectrum, not on UV compositeness.
@@ 5.3 Massive dark degrees of freedom: mapping to Yukawa envelopes
Setting $\mathcal{S}_{X}(1/2)=1$ gives a slightly conservative estimate that is convenient for quick re-interpretations of published $(\alpha,\lambda)$ curves (torsion balances, LLR, planetary dynamics, etc.). Applying the results from precision experiments at varying length scales, we can obtain continuous limits on the multiplicities of dark sectors, as is illustrated in fig. 1 below.
Including additional Yukawa poles when they enter the lever arm.
If the higher-derivative poles discussed in Sec. 2 descend into the laboratory window, their effect is incorporated by adding explicit Yukawa pieces to the static potential, rather than relying solely on the single-envelope mapping. We write
[…]
with $(\lambda_{k},\alpha_{k})$ the range and strength of the $k$-th Yukawa term. For the quadratic-curvature spin channels one may take $n_{\mathrm{Y}}=2$, with
$(\lambda_{0},\alpha_{0})\ \ \text{and}\ \ (\lambda_{2},\alpha_{2}),\qquad\lambda_{0,2}=\frac{1}{2M_{0,2}},$ (44)
mirroring the convention in Eq. (41). The data analysis then proceeds by profiling the likelihood (or $\chi^{2}$) over $(\alpha_{0},\alpha_{2})$ at fixed $(\lambda_{0},\lambda_{2})$, yielding one–sided bounds $\alpha_{0,2}^{95}(\lambda_{0,2})$ (single–pole case) or joint contours for two poles. Because Eq. (43) fits the sum of deviations from $1/r$, any additional pole within the probed range can only increase the total deviation unless it is finely tuned to cancel other pieces; our envelope construction avoids such tuned cancellations and thus remains conservative.
Figure 1: Degrees of freedom set by experimental constraints on $G$ and the Yukawa parameters $(\alpha,\lambda)$. The shaded region indicates the exclusion set by experimental limits at each distance scale $r_{\star}$.
@@ 6 Discussion and Conclusions
Note that our bounds are laboratory and model-independent: they constrain the present-day multiplicity of light states without assuming a cosmological population. Cosmology can provide stronger constraints when the hidden sector was populated and remained relativistic at BBN/CMB epochs, typically phrased as limits on $\Delta N_{\rm eff}$. However, low reheating temperatures, suppressed portals, late entropy injection, or purely gravitational/out-of-equilibrium production histories can substantially weaken cosmological sensitivity while leaving our reinterpretation intact. We therefore view the two approaches as complementary: CMB/BBN constrain abundances under specified histories, whereas precision gravity constrains the spectrum of light degrees of freedom irrespective of their relic density.
Our results have direct implications for several classes of theoretical models. Strongly coupled dark gauge sectors may contain multiple light states (e.g. PNGBs or glueballs) below the experimental lever arm; while each contributes negligibly, the cumulative effect scales with the number of such light states: any unexpectedly rich light spectrum is directly testable in the same, model-independent way. Our bounds provide quantitative guidance for the maximum complexity such sectors can achieve while remaining consistent with laboratory tests. Extra-dimensional theories often predict towers of Kaluza-Klein modes or large numbers of moduli fields, and the gravitational constraints complement existing bounds from fifth-force searches to help restrict the allowed parameter space of phenomenologically viable compactifications. Similar analyses could be extended to axion-like particles that couple to gravity through topological terms or induce effective modifications to gravitational couplings.
It is important to remark that Kaluza–Klein towers fit naturally into our framework provided one sums only over modes lighter than the probed scale. In the short-distance regime, massive KK gravitons generate a set of Yukawa deviations whose envelope is already captured by our mapping; the well-studied short-range gravity limits in ADD/RS scenarios can be rephrased in exactly this language. Our plots should thus be read as providing an efficient, model-agnostic translation between measured inverse-square-law tests and the cumulative effect of any tower of light modes.
We also note that for extremely large $N$, loop-induced higher-derivative poles can drift down toward the laboratory window. In that regime, our reinterpretation remains applicable and typically becomes more constraining, because such poles manifest as Yukawa deviations from $1/r$ that are bounded by the same precision data.